### v1

# Hummus: A Dataset of Humorous Multimodal Metaphor Use

Abstract
Metaphor and humor share a lot of common ground, and metaphor is one of the most common humorous mechanisms. This study focuses on the humorous capacity of multimodal metaphors, which has not received due attention in the community. We take inspiration from the Incongruity Theory of humor, the Conceptual Metaphor Theory, and the annotation scheme behind the VU Amsterdam Metaphor Corpus, and developed a novel annotation scheme for humorous multimodal metaphor use in image-caption pairs. We create the Hummus Dataset of Humorous Multimodal Metaphor Use, providing expert annotation on 1k image-caption pairs sampled from the New Yorker Caption Contest corpus. Using the dataset, we test state-of-the-art multimodal large language models (MLLMs) on their ability to detect and understand humorous multimodal metaphor use. Our experiments show that current MLLMs still struggle with processing humorous multimodal metaphors, particularly with regard to integrating visual and textual information. We release our dataset and code at github.com/xiaoyuisrain/humorous-multimodal-metaphor-use.
1 Introduction
Figure 1: Hummus metaphor sample, with annotation for image (bounding box and label), caption (<i></i>), conceptual metaphor, explanation of how the metaphor use contributes to humor, and additional figurative devices.
Conceptual Metaphor Theory (CMT; Lakoff and Johnson, 1980b) contends that the metaphors we use in language are based on cross-domain mappings in our mind; metaphor is essentially a cognitive process of understanding one thing in terms of another. For example, when one says Our marriage has gone off the track, one is using the love is a journey metaphor, conceptualizing love or relationship in terms of a journey. CMT marks a turning point in metaphor research, shifting focus towards how metaphors are represented or processed in the mind and the brain (see Holyoak and Stamenković, 2018 for a survey). There has also been an increased interest in nonverbal and multimodal manifestations of metaphor, such as metaphor in films, cartoons, adverts, and gestures Forceville (2015, 2017); Kappelhoff and Müller (2011); Tsakona (2009).
The influence of this cognitive turn in metaphor research has extended to the field of natural language processing. The VU Amsterdam Metaphor Corpus (VUA; Steen et al., 2010), which is created by cognitive linguists and employs an annotation scheme that closely follows the CMT, has inspired research on automatic detection of metaphors in text (see Tong et al., 2021 for a survey). There is also recent work on computational modelling of visual and multimodal metaphors in adverts, memes, and videos Alnajjar et al. (2022); Zhang et al. (2021, 2023); Xu et al. (2022).
Metaphor use serves many purposes in communication, including the delivery of humor Michelli et al. (2024); Attardo (2015). Metaphor and humor share a lot of common ground. The Incongruity Theory explains that humor arises from the perception of incongruity, something that violates one’s expectations Clark (1970); Morreall (2024). Similarly, metaphor researchers consider incongruity as a major contextual clue for identifying metaphors in text Cameron (2003); Steen et al. (2010). Metaphor and humor have even been suggested to share neural pathways in the brain Hellberg (2018), and visual metaphor is found to be one of the most common humorous mechanisms in cartoons Tsakona (2009).
This study is concerned with the humorous capacity of multimodal metaphors, and sheds light on how well multimodal large language models (MLLMs) understand humorous multimodal metaphor use. We created a Dataset of Humorous Multimodal Metaphor Use (Hummus), providing expert annotation on 1k image-caption pairs sampled from the New Yorker Caption Contest (CapCon) corpus Hessel et al. (2023) in terms of (1) whether the image-caption pair contains humorous multimodal metaphor use, (2) the conceptual metaphors involved, (3) image and text parts related to the metaphor use, (4) how the metaphor use contributes to humor, and (5) use of other figurative devices (e.g., idiom, irony, hyperbole). An example is provided in Figure 1. Our final dataset contains 558 metaphorical and 382 non-metaphorical items.
Based on the annotations, we created 6 tasks to test MLLMs’ ability to identify and understand humorous multimodal metaphor use. We tested four open-source and two closed source MLLMs: LLaVA-NeXT-110B/8B, Qwen2-VL-72B/7B-Instruct, GPT-4o, and GPT-4 Turbo, finding that the models struggle with distinguishing metaphorical and non-metaphorical items, as well as understanding the humorous multimodal metaphor use involved. Our ablation study and error analysis reveal that the models’ struggles are likely to be caused by difficulties in integrating visual and textual information.
2 Background
Metaphor.
CMT posits the existence of conceptual metaphors, mappings in the human mind between different conceptual domains Lakoff and Johnson (1980b). Metaphor as a figure of speech is manifestation of conceptual metaphors in language use. Conceptual metaphors also give rise to visual metaphors (e.g., representing an idea as a light bulb) and multimodal metaphors, which involve more than one mode of communication.
The image-caption pair in Figure 1 is a multimodal instantiation of the love is a journey metaphor. The caption explicitly compares sex life, which belongs to the love domain, with a train wreck, which can happen in a journey. The image depicts the love domain with a couple in bed, and the journey domain with a car crashing into the couple’s bedroom. The metaphor use involves both the linguistic and the visual mode, and thus qualifies as multimodal metaphor use.
Figurative language.
Metaphor, simile, personification, and zoomorphism are different types of figurative language, but they are all considered as metaphors under CMT. Similes are linguistic metaphors that use words such as like and as to highlight a cross-domain mapping (e.g., my love’s like a red, red rose); metaphor as a figure of speech uses categorization statements, making the cross-domain mapping more implicit (e.g., my love is a red, red rose). Personification is metaphor that compares something non-human to humans (e.g., pets as family members). Zoomorphism is metaphor with animals or a species of animal as the source domain (e.g., the roar of the ocean).
Idioms are sometimes called dead metaphors, expressions that have lost their metaphoricity over time. There is empirical evidence that people process idioms and metaphors differently Gibbs (1992); Desai (2022). We therefore regard simile, personification, and zoomorphism as metaphor use, but treat idiom as a different phenomenon.
Humor.
The Incongruity Theory of Humor can be traced back to Beattie (1779) and is the most influential theory for analyzing and explaining humor. The theory contends that laughter arises from the presence of incongruity, something that “violates our standard mental patterns and normal expectations Morreall (2024). Suls (1972) proposes a two-stage model for humor appreciation: perception of incongruity and resolution of incongruity. For example, the cartoon image in Figure 1 is incongruous as car crash normally does not happen in a bedroom. The caption resolves the incongruity by attributing the car crash to the couple’s disastrous relationship, and humor arises from this witty, unexpected attribution.
3 Related Works
Linguistic metaphor.
A large proportion of research on computational modelling of metaphor is restricted to linguistic metaphors, and automatic metaphor detection in particular Li et al. (2023); Maudslay and Teufel (2022); Shutova (2015); Tong et al. (2021); Wang et al. (2023); Zhang and Liu (2022). Most metaphor detection studies employ the VU Amsterdam Metaphor Corpus (VUA; Steen et al., 2010), a four-million word corpus where every word is annotated in terms of whether its use is metaphorical.
[…]
Visual and multimodal metaphor.
Automatic identification of text-based metaphors in videos Alnajjar et al. (2022) and figurative language use in memes Liu et al. (2022a) has been tackled. For metaphor understanding, the V-FLUTE dataset Saakyan et al. (2024) provides automatically generated explanations for visual metaphors that are automatically generated from linguistic metaphors. Other datasets provide detailed annotations for monomodal and multimodal metaphors in social media and adverts Zhang et al. (2021, 2023) or memes Xu et al. (2022). Their annotations concern metaphor occurrence, modality (whether the metaphor is text-based, image-based, or multimodal), target domain, source domain, intent, and sentiment category.
Akula et al. (2023) annotate visual metaphors in adverts and introduce MetaCLUE, a set of five visual metaphor processing tasks: classification (whether a given image contains metaphor), localization (identifying image regions), understanding (interpreting metaphors in the form “<target> is as <property> as <source>”), and generation (generating visual metaphors from text prompt).
Humor and metaphor.
Hessel et al. (2023) propose a set of humor understanding tasks based on their New Yorker Caption Contest (CapCon) corpus. The corpus contains earlier releases Jain et al. (2020); Radev et al. (2016); Shahaf et al. (2015) of The New Yorker cartoons and captions. It also provides manually created explanations for 651 image-caption pairs.
Chang et al. (2024) annotate a subset of the CapCon corpus and propose NYK-MS, a benchmark for metaphor and sarcasm understanding. While our work also uses the CapCon corpus, our contributions are different in the following ways: (1) Every metaphorical item in Hummus qualifies as multimodal metaphor use, and its multimodality is reflected in the annotations; NYK-MS, on the other hand, is primarily concerned with metaphorical words in captions. (2) We annotate the conceptual metaphors underlying the identified multimodal metaphor use; NYK-MS lacks this type of annotation. (3) We provide bounding box annotations for image areas related to the identified metaphor use, which is also absent in NYK-MS. (4) We explain our metaphorical items in relation to the humorous effect of the image-caption pairs, whereas NYK-MS does not consider the interplay between metaphor and humor. (5) Hummus is annotated by a linguist specializing in metaphor research while the ground truth of NYK-MS is based on GPT-4V generations.
4 Dataset Creation
Hummus is built upon the image-caption pairs in the CapCon corpus Hessel et al. (2023). Each week, The New Yorker publishes a captionless cartoon and receives caption submissions from its readers. The CapCon corpus includes 2578 funny caption submissions for 679 high-quality cartoons of The New Yorker. We randomly sampled 251 cartoons from the corpus, for which a total of 1000 image-caption pairs are available. We designed an annotation scheme informed by theories of humor understanding and metaphor use, and a volunteer linguist with expertise in metaphor annotated the image-caption pairs. Our annotation scheme can be summarized into two stages: humorous metaphor identification and detailed metaphor annotation, which we explain below.
4.1 Humorous metaphor identification
Humorous metaphor use is identified in two steps: humor understanding and metaphor identification. Image-caption pairs are tagged “Yes”, “No”, “WIDLII (While In Doubt, Leave It In)”, or “Discard” at this stage. Given an image-caption pair, the annotator first employs the incongruity-resolution approach to understand the humor: They note down all possible incongruities in the image and see how the caption resolves those incongruities. If the annotator fails to understand the humor (incongruities remain unresolved), the item is tagged as “Discard”.
If the annotator understands the humor, they proceed to determine whether the humor involves metaphor use—whether it can be attributed to any cross-domain mapping, or a process of conceptualizing/depicting one thing in terms of another. If the answer is a definite yes or no, the item is marked as “Yes” or “No” respectively. If the annotator is uncertain, the item is marked as “WIDLII” instead. We adopt this strategy from the annotation scheme behind the VUA corpus Steen et al. (2010). Both “Yes” and “WIDLII” items participate in the subsequent stage of detailed metaphor annotation; this allows us to avoid mistakenly marking an interesting or implicit case of metaphor use as non-metaphorical.
4.2 Detailed metaphor annotation
Figure 2: Metaphor sample for which multiple conceptual metaphors are annotated: pub is a coal mine; humans are animals.
Image-caption pairs that are tagged as “Yes” or “WIDLII” in terms of the involvement of humorous metaphor use are further annotated in three aspects: the involved conceptual metaphor, how the conceptual metaphor is reflected in the image (highlighting the relevant objects) and caption (highlighting relevant words), and an explanation of how the metaphor use contributes to the humorous effect of the image-caption pair. WIDLII items could be re-labelled as No or Discard items; an explanation is provided if a WIDLII item is eventually Discarded.
Conceptual metaphor.
Following the tradition of previous research Lakoff and Johnson (1980a); Lakoff et al. (1991), the underlying conceptual metaphors are annotated in target domain is source domain format.
An image-caption pair could employ multiple conceptual metaphors to achieve humorous effects. The annotator is thus asked to be as inclusive as possible, specifying all conceptual metaphors they can identify. Take the image-caption pair in Figure 2 as an example. The image shows two people having drinks in a pub. One of them wears a hazmat suit, which is unheard-of in pubs. The caption justifies this strange clothing choice (thus resolving the incongruity) by revealing that the person considers himself in a coal mine. There is thus a metaphorical comparison between pub and coal mine. The caption also explicitly compares the man in an ordinary suit to a canary in a coal mine (in its literal sense), an instantiation of the humans are animals metaphor.
Image and caption annotation.
The annotator also marks out image and text fragments that are related to the annotated conceptual metaphor(s). Metaphor-related image areas are annotated using both bounding boxes and texts (a word or phrase that tells which part of the image is highlighted); each bounding box has a corresponding textual description. Metaphor-related words or phrases are surrounded by <i></i> tags.
A conceptual metaphor is a mapping between two conceptual domains. To annotate the image and the caption thus requires the annotator to determine how the two domains are reflected in the two modalities. Usually, representations of the two domains are unbalanced: The image should have a recognizable setting that represents the target or the source domain of the metaphor, while a small part of the image (e.g., a particular object) points to the other domain, thus creating incongruity and cross-domain mapping. Similarly, the context of the caption can be assigned to the target or the source domain, while a particular word or phrase suggests the involvement of the other domain. The annotator’s job is therefore to mark out the less represented domain in the image and the caption respectively.
Let us return to the image-caption pair in Figure 1, which illustrates the love is a journey metaphor. The image predominantly belongs to the love domain, with a bedroom as the setting and the couple sitting in bed occupying more than half of the image. The car crashing into the room on the left side of the image evokes the journey domain, creating incongruity and encouraging cross-domain mapping. While the cracks in the ceiling, the damaged door and the mess on the ground come along with the car crash, they are visual representations of the result of the cross-domain mapping, as opposed to belonging to the source domain, journey, itself. The car is thus annotated as the metaphor-related fragment in the image. The caption talks about the couple’s sex life and compares it explicitly with a train wreck. The phrase train wreck is the “incongruous” part of the caption, evidence of the journey domain in the context of the love domain. It is therefore marked out as metaphor-related.
Explanation.
For each image-caption pair that contains humorous multimodal metaphor use, we also provide a short explanation about how the metaphor use contributes to the humor. This is different from the explanations provided in the CapCon corpus: While the CapCon corpus focuses on humor understanding and only explicitly mentions metaphor use for two image-caption pairs, our explanations, as exemplified in Figure 1, focus specifically on the interplay between humor and metaphor use.
Other figurative devices.
We also provide annotation of the use of other figurative devices, such as pun, idiom, irony, hyperbole, for both metaphorical and non-metaphorical image-caption pairs. This part of the annotation provides insight into the relation between humor and the use figurative expressions in general, as well as co-occurrence of metaphor and other figures of speech in delivering humor. We do not use these data to test MLLMs in this study, but still include them in the dataset, as they could be of interest for future research on figurative language.
5 Data Analysis
(a) animals are humans.
(b) animals/humans are humans/animals.
(c) Go through the roof.
Figure 3: Example of (a) unidirectional and (b) bidirectional metaphorical mappings between animals and humans, and example of (c) metaphorically used idiom in our dataset.
Hummus provides annotations for 1,000 image-caption pairs, including 558 items that contain humorous multimodal metaphor use (331 “Yes” and 227 “WIDLII”), 382 “No” items, and 60 items marked as “Discard”.
Multimodality of the metaphor samples.
Are the metaphors annotated in our dataset identifiable from the captions alone? To answer this question, we employ FrameBERT, an automatic textual metaphor detector proposed by Li et al. (2023). The model reaches an F1 score of 0.73 when tested on the VUA-20 dataset Leong et al. (2020). We consider it as a “Yes” answer if the model predicts that a given caption contains a content-word real metaphor or borderline metaphor. Compared with ground truth, the model reaches an accuracy of 0.53 and a F1 score of 0.54, both of which are close to a random guess, and also lower than the model’s performance on detecting linguistic metaphors. This can be considered as an indicator of the multimodal nature of the metaphors annotated in our dataset.
Conceptual metaphor.
Around 390 conceptual metaphors are identified in the dataset. The most frequently occurring ones are animals are humans (23% metaphor samples) and humans are animals (8%). In addition to comparison with animals, the dataset also includes mappings from humans as a source domain to other target domains such as object, god, and aliens. Personification thus appears to be the most frequently used type of metaphor in these humorous image-caption pairs (51%).
Directionality of metaphors.
A crucial and inevitable step in identifying any (conceptual) metaphor is determining the direction of the metaphorical mapping—in other words, which of the two conceptual domains at play is the target domain, and which is the source domain. Metaphorical mappings are unidirectional: Properties and relations of the source domain are projected onto the target domain, not the other way around. One cannot reverse the direction of a mapping without creating an entirely different metaphor. Consider the image-caption pair in Figure 3(a), for example. It is clear that the joke is based on a metaphor with animals as the target domain and humans as the source domain: A hippo is given human characteristics—it uses telephones and gets angry when someone calls it at an inconvenient time.
This rule of unidirectionality applies to “prototypical metaphors of all kinds and occurring in all media” Forceville (2002) while exceptions also exist Carroll (1994); Forceville (1995). Hummus also demonstrates some exceptions to the rule, especially with regard to the humans and animals domains. The image-caption pair in Figure 3(b), for example, shows a cat having a meeting with some other animals in a modern conference room. The cat expresses its wish to keep the meeting brief, so that it could go back to staring out the window. The fact that the animals are sitting in a conference room and engaged in a meeting can be considered as employing the animals are humans metaphor. The humor of the image-caption pair thus lies in the cat attaching ample importance to staring out the window despite its position in the company; it also invites the reader to wonder what kind of business this company might be running.
On the other hand, one can also interpret the joke as based on the metaphor humans are animals—people working in a company are represented as animals in the image. By representing the person in a higher position as a cat, the metaphor satirically emphasizes the absurdity of a common scenario in society: A person has an important position in a company while all they care about is something as unproductive as staring out the window. For this image-caption pair, therefore, it is difficult to distinguish humans and animals in terms of target and source domains, unless we ask the caption writer directly; but it could also happen that the caption writer intended the metaphor to be bidirectional in the first place. Nonetheless, the double interpretation adds to the depth of such image-caption pairs, making them particularly interesting cases for both metaphor analysis and model evaluation.
Co-occurrence with other figurative devices.
More than half (65%) of the metaphor samples in Hummus feature metaphor use co-occurring with the use of other figurative devices (excluding personification and zoomorphism, which are considered metaphors in this study). The most frequently used ones include pun (27%), exaggeration (10%), and satire (10%).
We also find a small percentage of metaphor samples (5%) where the humorous metaphor use concerns an idiom—recall that idioms are usually considered dead metaphors (Section 2). Consider the image-caption pair in Figure 3(c). When we look at the caption alone, its use of the idiom go through the roof is non-metaphorical: One understands it refers to a high cholesterol level without visualizing cholesterol actually going through the roof. When the caption is combined with the image, however, the metaphorical mapping between amount and height is resurrected, and it is precisely the resurrection of the dead metaphor that brings out humor.
6 Model Evaluation
Task Instructions Evaluation
Classification Does the humor of the given image-and-caption combination involve metaphor use? Answer the question with Yes or No. Multiclass F1 score
Naming The humor of the given image-and-caption combination involves metaphor use. Which conceptual metaphor is used? Answer the question in "TARGET DOMAIN IS SOURCE DOMAIN" format (e.g., "LOVE IS A JOURNEY"). Sentence similarity
ImageBbox The humor […] involves metaphor use. Which object in the image is related to the metaphor? Answer with its label and normalized bounding box coordinates in "label: [top, left, height, width]" format. IoU, precision, recall
ImageLabel The humor […] involves metaphor use. Which object in the image is related to the metaphor? Answer the question with a single word. Sentence similarity
CaptionHL The humor […] involves metaphor use. Which part of the caption is related to the metaphor? Surround it with a pair of <i></i> tag. Jaccard index, precision, recall
Explanation How does metaphor use contribute to the humor of the given image-and-caption combination? Explain in no more than 30 words. ROUGE-1, ROUGE-2
[…]
We design six tasks for humorous multimodal metaphor processing: Classification, Naming, ImageBbox, ImageLabel, CaptionHL, and Explanation. Table 1 provides an overview of the instructions and evaluation metrics for each task. The Classification task includes all items in the test set—that is, the image-caption pair is tagged “Yes”, “No”, or “WIDLII” in terms of whether or not it involves humorous multimodal metaphor use. The “No” items are considered negative cases in the Classification task; the “Yes” and “WIDLII” items positive, as our dataset provides full metaphor annotation for both categories. All other tasks only involve the positive cases.
The Naming task and the ImageLabel task employ LaBSE Feng et al. (2022) as the evaluator. Our choice is based on a pilot Naming test that involves the first 100 items in our test set and two models: GPT-4o and GPT-4 Turbo. We calculate cosine similarity scores between the model outputs and ground truth using a variety of SBERT models. We choose LaBSE as its predictions are the closest to human judgement.
Model Classification Naming Explanation
Pos Neg Avg ROUGE-1 ROUGE-2
GPT-4 Omni 0.70 0.39 0.55 0.47 (0.22) 0.23 (0.08) 0.03 (0.04)
Turbo 0.64 0.47 0.56 0.49 (0.21) 0.22 (0.08) 0.03 (0.04)
LLaVA-NeXT 110B 0.69 0.36 0.53 0.43 (0.18) 0.27 (0.08) 0.05 (0.04)
8B 0.74 0.00 0.37 0.46 (0.18) 0.26 (0.08) 0.04 (0.04)
Qwen2-VL 72B 0.64 0.43 0.54 0.44 (0.18) 0.25 (0.08) 0.04 (0.04)
7B 0.61 0.46 0.53 0.40 (0.18) 0.26 (0.08) 0.04 (0.04)
Random 0.54 0.45 0.50
Table 2: Classification, Naming, and Explanation benchmarks. Standard deviations are in parenthesis. GPT-4 Turbo has a success rate less than 100% in the Naming task (96%).
Model ImageBbox ImageLabel Caption
P R IoU P R Jaccard
GPT-4 Omni 0.40 (0.34) 0.43 (0.34) 0.25 (0.22) 0.63 (0.29) 0.47 (0.36) 0.85 (0.32) 0.42 (0.32)
Turbo 0.38 (0.37) 0.32 (0.34) 0.19 (0.21) 0.56 (0.27) 0.54 (0.39) 0.80 (0.37) 0.47 (0.36)
LLaVA-NeXT 110B 0.17 (0.20) 0.50 (0.38) 0.14 (0.16) 0.58 (0.27) 0.42 (0.38) 0.93 (0.24) 0.40 (0.37)
8B 0.17 (0.20) 0.64 (0.39) 0.15 (0.16) 0.58 (0.27) 0.33 (0.33) 0.95 (0.20) 0.31 (0.30)
Qwen2-VL 72B 0.17 (0.20) 0.40 (0.35) 0.13 (0.14) 0.59 (0.28) 0.40 (0.37) 0.89 (0.29) 0.36 (0.34)
7B 0.17 (0.21) 0.40 (0.34) 0.13 (0.14) 0.57 (0.27) 0.36 (0.35) 0.94 (0.22) 0.34 (0.32)
Table 3: ImageBbox, ImageLabel, and CaptionHL benchmarks in mean (SD) format. A few tests result in success rates less than 100%: no less than 98% in the ImageBbox task; 97% for LLaVA-NeXT-8B, 89% for LLaVA-NeXT-110B, and 41% for Qwen2-VL-7B in the CaptionHL task.
We evaluate six models, including four state-of-the-art MLLMs, GPT-4 Omni (GPT-4o, gpt-4o-2024-05-13), GPT-4 Turbo (gpt-4-turbo-2024-04-09), LLaVA-NeXT-110B Liu et al. (2023a, b, 2024), and Qwen2-VL-72B-Instruct Wang et al. (2024), and two smaller, open-source models, LLaVA-NeXT-8B, Qwen2-VL-7B-Instruct. The smaller models require less computing resources and can be more versatile than their larger counterparts in certain use cases (e.g., for fine-tuning).
6.1 Benchmark results
Table 2 and 3 show results of testing the six models on the six tasks presented in Table 1. We also report success rates, which measure whether the models provide meaningful answers for evaluation (e.g., if a model merely repeats the instructions, it fails to provide a meaningful answer).
The random baseline for the Classification task is calculated by randomly choosing between a “Yes” and a “No” answer for each item, and averaging the results of 100 iterations. For ImageBbox, IoU scores above 0.5 are considered a match between model prediction and ground truth. For Naming, ImageLabel, and Explanation, we decide thresholds for good answers by manually examining model answers in various score ranges. The threshold for Naming and ImageLabel is set at 0.6. For Explanation, model answers reaching a ROUGE-1 score of 0.35 and a ROUGE-2 score of 0.087 are considered acceptable.
How well do the models identify humorous multimodal metaphor use?
All six models are prone to classifying the image-caption pairs as metaphorical. While they achieve F1 scores higher than the random baseline in identifying positive cases, even the highest F1 score for the negative category (0.47 by GPT-4 Turbo) is merely around the random baseline (0.45).
How well do the models identify the underlying conceptual metaphors?
All models struggle with identifying the underlying conceptual metaphors. Only a small proportion of model predictions (21%) reachs the threshold of 0.6 in the Naming task.
How well do the models localize humorous multimodal metaphor use in image and caption?
The models are much better at labelling metaphor-related image areas than providing their coordinates: Around 40% model answers reach the 0.6 threshold in the ImageLabel task, whereas the highest mean IoU score in the ImageBbox task is 0.25 (GPT-4o), which suggests barely any overlap with the ground truth.
When highlighting metaphor-related text fragments, the models are able to cover a large proportion of ground truth annotations, but their answers also tend to be longer, including text fragments that are not related to the humorous metaphor use.
How well do the model explain humorous multimodal metaphor use?
The Explanation task proves to be difficult for the models. Merely 5% of model predictions reach our threshold of acceptable explanations.
6.2 Prompt engineering
Model Positive Negative Average
GPT-4o 0.71 (0.01) 0.43 (0.02) 0.57 (0.01)
Qwen2-VL-72B 0.61 (0.06) 0.46 (0.07) 0.53 (0.01)
Qwen2-VL-7B 0.66 (0.06) 0.36 (0.11) 0.51 (0.03)
Table 4: Mean F1 scores using 3 different prompts for the Classification task. Qwen2 stands for Qwen2-VL-Instruct. Success rate is always 100%.
Figure 4: Model performance in multimodal versus monomodal experiments: average F1 score for Classification, sentence similarity score for Naming and ImageLabel, Jaccard index score for CaptionHL, ROUGE-1 for Explanation. Success rate is 100% except for Qwen2-VL-7B-Instruct in the monomodal CaptionHL task (97%).
We experiment with other ways to formulate the Classification task, to see whether the high probability of “Yes” answers is associated with the prompt we use. Using the first 100 items in the test set, we run a pilot study that tests a wide range of prompts on one of the worst performing models, Qwen2-VL-7B-Instruct. Instead of asking the model to reply with “Yes” or “No”, these prompts require different ways to label the given image-caption pair, such as “Metaphorical/Non-metaphorical”, “True/False”, “A/B”.
We determine the top-3 prompts that result in the highest average F1 scores in the pilot study. These prompts ask the models to reply with (1) “Yes (i.e., metaphor use is involved) or No (i.e., metaphor use is not involved)”, (2) “No or Yes”, and (3) “A or B”, respectively. For the third prompt, the two options correspond to whether or not the given image-caption pair involves humorous metaphor use (i.e., “involves” or “does not involve”), and the order of the two options is randomised for each test item.
We test these three prompts on three models: GPT-4o, and the two Qwen2-VL models. GPT-4o is one of the best-forming models over all tasks, and its success rate is always 100%. The two Qwen2-VL models represent the best and the worst open-source models. The 72B model has 100% success rate on all tasks. The 7B model is the most unstable, achieving a 41% success rate on the CaptionHL task.
As shown in Table 4, the two Qwen2-VL models, especially the smaller one, are sensitive to different prompts. The performance of GPT-4o, on the other hand, remains stable over different prompts. The experiment thus proves the reliability of our Classification benchmark results.
6.3 Ablation study
To examine the models’ processing of multimodal input in our tasks, we design an ablation study that replaces image input with textual descriptions of the images, making the input data purely textual. We use the image_description data provided in the CapCon corpus, which are short, literal descriptions of the scene. For example, the description for the image in Figure 1 is as follows: “A man and woman are in bed together under the covers. They are looking towards the bedroom door when a car has crashed into their home. They don’t seem too upset about the situation.”
We rerun the six tasks on the same three models for prompt engineering: GPT-4o and the two Qwen2-VL models. For Classification, we use the prompt that results in the highest average F1 score for most models in prompt engineering: It asks the models to answer the question with “No or Yes” instead of “Yes or No”.
As shown in Figure 4, there is not much difference between the models’ performance in the multimodal and monomodal experiments, except for GPT-4o in the Classification task. The comparison indicates that the multimodal input data is not adequately utilized by the models.
6.4 Error Analysis
Figure 5: Ground truth versus LLaVA-NeXT-110B predictions for the given image-caption pair. Explanations are shortened. For CaptionHL, the model outputs the entire caption as answer.
Figure 6: All models except GPT-4o predict life is a vacuum for this item in the Naming task.
A primary reason for incorrect answers is the models’ inability to integrate visual and textual information into a coherent story. For example, the humor of the image-caption pair in Figure 5 is based on personification of the sharks: They communicate in human language and can experience and express embarrassment. Subsequently, the person in the image is compared metaphorically to an object, referred to as it in the caption.
LLaVA-NeXT-110B’s answers indicate that the model assumes the caption is uttered by the person in the image, thus overlooking the animals are humans metaphor and the subsequent a person is an object metaphor. This example also shows the importance of metaphor processing in humor understanding: One cannot understand the humor of this image-caption pair without recognizing that the sharks are personified.
On the other hand, the models can usually identify common metaphorical or idiomatic expressions in the caption, although it does not guarantee adequate understanding of humorous multimodal metaphor use. For example, the cartoon in Figure 6 depicts an alien spaceship as a vacuum cleaner. The humor comes from the pun on vacuum in the caption: It refers both to aliens physically existing in a vacuum cleaner, as depicted in the image, and life existing in a vacuum in an idiomatic sense. The ground truth annotation of the conceptual metaphor is alien spaceship is a vacuum cleaner, because it is the base of the pun on vacuum in the caption. The MLLMs acknowledge the metaphoricity of the expression life can’t exist in a vacuum, but their prediction in the Naming task, life is a vacuum, indicates that they do not relate it to the vacuum cleaner spaceship in the image, thus failing to properly understand the humorous multimodal metaphor use.
Our analysis indicates that the models’ performance can potentially be improved when they are instructed more explicitly to combine image and caption information. For example, one can use chain-of-thought prompting to instruct an MLLM to mimic the human annotation process: first process the image, and then integrate it with the caption, before proceeding to metaphor identification and understanding.
7 Conclusion
This study releases a dataset that provides expert annotation on humorous multimodal metaphor use. Using the dataset, we benchmark popular and state-of-the-art MLLMs on their capabilities to identify and understand humorous multimodal metaphor use. Our experiments show that current MLLMs struggle with processing humorous multimodal metaphor use, especially with regard to integrating visual and textual information.
Limitations and future directions.
We hope this study will encourage more research on MLLMs’ capabilities to process humorous multimodal metaphors. Our dataset is annotated by a single expert. While this ensures efficient, high-quality, and systematic labelling, it may also introduce a certain level of subjectivity to the dataset. We envisage future work exploring alternative interpretations of the image-caption pairs and how MLLMs align with such individuality.
We also acknowledge that the latest reasoning models such as DeepSeek R1 and OpenAI o1 are not included in this study. They are released after the experiments were conducted and do not necessarily have multimodal capabilities (e.g., DeepSeek R1).
8 Ethical Considerations
We access the GPT-4 models through OpenAI API. The open-source models (LLaVA-NeXT-8/110B and Qwen2-VL-7/72B) as well as the CapCon corpus are accessed through HuggingFace. The CapCon corpus has CC-BY-4.0 license. Our dataset will be freely accessible on GitHub.
Our dataset includes jokes that could be considered offensive, and certain jokes may be inappropriate for a younger audience. These data remain in our dataset as they can be valuable for future research.

### v2

@@
 Our final dataset contains 558 metaphorical and 382 non-metaphorical items.
-Based on the annotations, we created 6 tasks to test MLLMs’ ability to identify and understand humorous multimodal metaphor use.
+Based on the annotations, we created 6 tasks to test MLLMs’ ability to identify and understand humorous multimodal metaphor use: Classification, Naming (of conceptual metaphors), ImageBbox, ImageLabel, CaptionHL, and Explanation.
 We tested four open-source and two closed source MLLMs: LLaVA-NeXT-110B/8B, Qwen2-VL-72B/7B-Instruct, GPT-4o, and GPT-4 Turbo, finding that the models struggle with distinguishing metaphorical and non-metaphorical items, as well as understanding the humorous multimodal metaphor use involved.
@@
 Metaphor.
-CMT posits the existence of conceptual metaphors, mappings in the human mind between different conceptual domains Lakoff and Johnson (1980b).
+The Conceptual Metaphor Theory posits the existence of conceptual metaphors, mappings in the human mind between different conceptual domains Lakoff and Johnson (1980b).
 Metaphor as a figure of speech is manifestation of conceptual metaphors in language use.
@@
 Most metaphor detection studies employ the VU Amsterdam Metaphor Corpus (VUA; Steen et al., 2010), a four-million word corpus where every word is annotated in terms of whether its use is metaphorical.
-[…]
+Nevertheless, recent years have seen an increased interest in metaphor understanding.
+More paraphrase datasets are released Joseph et al.
+(2023); Tong et al.
+(2024), following earlier works that treat metaphor understanding as a paraphrasing task Bizzoni and Lappin (2018); Shutova (2010).
+Other studies frame the issue as an inference task, providing entailed and non-entailed Com s , a et al.
+(2022); Stowe et al.
+(2022) or contradictory statements Chakrabarty et al.
+(2022); Liu et al.
+(2022b) for metaphorical sentences.
+A framework for intentions behind metaphor use has also been proposed Michelli et al.
+(2024).
 Visual and multimodal metaphor.
@@
 (2023).
-Each week, The New Yorker publishes a captionless cartoon and receives caption submissions from its readers.
-The CapCon corpus includes 2578 funny caption submissions for 679 high-quality cartoons of The New Yorker.
+Each week, The New Yorker publishes a captionless cartoon and receives caption submissions from readers.
+The CapCon corpus includes 2578 funny caption submissions for 679 cartoons of The New Yorker.
 We randomly sampled 251 cartoons from the corpus, for which a total of 1000 image-caption pairs are available.
-We designed an annotation scheme informed by theories of humor understanding and metaphor use, and a volunteer linguist with expertise in metaphor annotated the image-caption pairs.
-Our annotation scheme can be summarized into two stages: humorous metaphor identification and detailed metaphor annotation, which we explain below.
+Each image-caption pair is a unique instance in terms of its humor and possible metaphor use.
+We designed an annotation scheme informed by theories of humor understanding and metaphor use, and manually annotated the image-caption pairs on Labelbox.
+Our annotation scheme can be summarized into two stages: (1) humorous metaphor identification, which takes less than a minute per item, and (2) detailed metaphor annotation, which takes $\sim$15 minutes per item.
 4.1 Humorous metaphor identification
@@
 We do not use these data to test MLLMs in this study, but still include them in the dataset, as they could be of interest for future research on figurative language.
-5 Data Analysis
+4.3 Inter-annotator agreement
+A second expert annotated 300 image-caption pairs randomly sampled from the 1000.
+Like the first annotator, the second annotator is a linguist with knowledge of CMT and experience in linguistic metaphor annotation.
+To ensure timely completion of the task, we left out explanation and figurative device annotation and asked the second expert to (1) classify the items in terms of humorous metaphor use (Yes/No/WIDLII/Discard), (2) identify the conceptual metaphors used in the Yes/WIDLII items, and (3) annotate their image (bounding box) and caption.
+To facilitate comparison with model performance, we use the same metrics for inter-annotator agreement calculation and model evaluation: multicalss F1 score for humorous metaphor identification, cosine similarity for conceptual metaphor identification, Intersection over Union (IoU) for image annotation, and Jaccard index for caption annotation.
+The co-annotation process started with us introducing the second annotator to our annotation scheme.
+We provided them with the written annotation guidelines and verbally walked them through every step of the scheme (excluding explanation and figurative device annotation), exemplifying both metaphorical and non-metaphorical instances.
+After the first meeting, the second annotator was assigned the first 50 items to complete independently.
+Upon finishing the first 50 items, the two annotators inspected their disagreements together and discussed necessary amendments to the original annotation guidelines.
+After the second meeting, we updated the guidelines and let the second annotator finish the remaining 250 items independently.
+The annotation was completed in 2 months.
+With sufficient familiarity with the procedure, annotation for a metaphorical item can be completed in 7 minutes.
+The guidelines are provided in Appendix A.
+The final agreement scores were calculated without further discussion of disagreements: 0.73 average F1 score for humorous metaphor identification, with 0.75 for metaphorical (Yes/WIDLII) items and 0.71 for non-metaphorical (No) items; 0.63 mean similarity score for conceptual metaphor identification ($SD=0.20$); 0.73 mean IoU score for image annotation ($SD=0.32$); 0.65 mean Jaccard index for caption annotation ($SD=0.41$).
+Figure 3: Conceptual metaphors identified in our dataset. Font size corresponds to frequency of occurrence.
 (a) animals are humans.
@@
 (c) Go through the roof.
-Figure 3: Example of (a) unidirectional and (b) bidirectional metaphorical mappings between animals and humans, and example of (c) metaphorically used idiom in our dataset.
+Figure 4: Example of (a) unidirectional and (b) bidirectional metaphorical mappings between animals and humans, and example of (c) metaphorically used idiom (go through the roof) in our dataset.
+5 Data Analysis
 Hummus provides annotations for 1,000 image-caption pairs, including 558 items that contain humorous multimodal metaphor use (331 “Yes” and 227 “WIDLII”), 382 “No” items, and 60 items marked as “Discard”.
@@
 Around 390 conceptual metaphors are identified in the dataset.
-The most frequently occurring ones are animals are humans (23% metaphor samples) and humans are animals (8%).
+As shown in Figure 3, The most frequently occurring ones are animals are humans (23% metaphor samples) and humans are animals (8%).
 In addition to comparison with animals, the dataset also includes mappings from humans as a source domain to other target domains such as object, god, and aliens.
 Personification thus appears to be the most frequently used type of metaphor in these humorous image-caption pairs (51%).
+Task Instructions Evaluation
+Classification Does the humor of the given image-and-caption combination involve metaphor use? Answer the question with Yes or No. Multiclass F1 score
+Naming The humor of the given image-and-caption combination involves metaphor use.
+Which conceptual metaphor is used?
+Answer the question in "TARGET DOMAIN IS SOURCE DOMAIN" format (e.g., "LOVE IS A JOURNEY").
+Sentence similarity
+ImageBbox The humor […] involves metaphor use.
+Which object in the image is related to the metaphor?
+Answer with its label and normalized bounding box coordinates in "label: [top, left, height, width]" format.
+IoU, precision, recall
+ImageLabel The humor […] involves metaphor use. Which object in the image is related to the metaphor? Answer the question with a single word. Sentence similarity
+CaptionHL The humor […] involves metaphor use. Which part of the caption is related to the metaphor? Surround it with a pair of <i></i> tag. Jaccard index, precision, recall
+Explanation How does metaphor use contribute to the humor of the given image-and-caption combination? Explain in no more than 30 words. ROUGE-1, ROUGE-2
+[…]
 Directionality of metaphors.
@@
 6 Model Evaluation
-Task Instructions Evaluation
-Classification Does the humor of the given image-and-caption combination involve metaphor use? Answer the question with Yes or No. Multiclass F1 score
-Naming The humor of the given image-and-caption combination involves metaphor use.
-Which conceptual metaphor is used?
-Answer the question in "TARGET DOMAIN IS SOURCE DOMAIN" format (e.g., "LOVE IS A JOURNEY").
-Sentence similarity
-ImageBbox The humor […] involves metaphor use.
-Which object in the image is related to the metaphor?
-Answer with its label and normalized bounding box coordinates in "label: [top, left, height, width]" format.
-IoU, precision, recall
-ImageLabel The humor […] involves metaphor use. Which object in the image is related to the metaphor? Answer the question with a single word. Sentence similarity
-CaptionHL The humor […] involves metaphor use. Which part of the caption is related to the metaphor? Surround it with a pair of <i></i> tag. Jaccard index, precision, recall
-Explanation How does metaphor use contribute to the humor of the given image-and-caption combination? Explain in no more than 30 words. ROUGE-1, ROUGE-2
-[…]
-We design six tasks for humorous multimodal metaphor processing: Classification, Naming, ImageBbox, ImageLabel, CaptionHL, and Explanation.
-Table 1 provides an overview of the instructions and evaluation metrics for each task.
-The Classification task includes all items in the test set—that is, the image-caption pair is tagged “Yes”, “No”, or “WIDLII” in terms of whether or not it involves humorous multimodal metaphor use.
-The “No” items are considered negative cases in the Classification task; the “Yes” and “WIDLII” items positive, as our dataset provides full metaphor annotation for both categories.
-All other tasks only involve the positive cases.
-The Naming task and the ImageLabel task employ LaBSE Feng et al.
-(2022) as the evaluator.
-Our choice is based on a pilot Naming test that involves the first 100 items in our test set and two models: GPT-4o and GPT-4 Turbo.
-We calculate cosine similarity scores between the model outputs and ground truth using a variety of SBERT models.
-We choose LaBSE as its predictions are the closest to human judgement.
 Model Classification Naming Explanation
@@
 Random 0.54 0.45 0.50
-Table 2: Classification, Naming, and Explanation benchmarks. Standard deviations are in parenthesis. GPT-4 Turbo has a success rate less than 100% in the Naming task (96%).
+Table 2: Classification, Naming, and Explanation benchmarks. Standard deviations are in parenthesis. Success rate is 100% except GPT-4 Turbo in the Naming task (96%).
 Model ImageBbox ImageLabel Caption
@@
 A few tests result in success rates less than 100%: no less than 98% in the ImageBbox task; 97% for LLaVA-NeXT-8B, 89% for LLaVA-NeXT-110B, and 41% for Qwen2-VL-7B in the CaptionHL task.
+We design six tasks for humorous multimodal metaphor processing: Classification, Naming, ImageBbox, ImageLabel, CaptionHL, and Explanation.
+Table 1 provides an overview of the instructions and evaluation metrics for each task.
+The Classification task includes all items in the test set—that is, the image-caption pair is tagged “Yes”, “No”, or “WIDLII” in terms of whether or not it involves humorous multimodal metaphor use.
+The “No” items are considered negative cases in the Classification task; the “Yes” and “WIDLII” items positive, as our dataset provides full metaphor annotation for both categories.
+All other tasks only involve the positive cases.
+The Naming task and the ImageLabel task employ LaBSE Feng et al.
+(2022) as the evaluator.
+Our choice is based on a pilot Naming test that involves the first 100 items in our test set and two models: GPT-4o and GPT-4 Turbo.
+We calculate cosine similarity scores between the model outputs and ground truth using a variety of SBERT models.
+We choose LaBSE as its predictions align the most closely to human judgement of good and bad answers.
 We evaluate six models, including four state-of-the-art MLLMs, GPT-4 Omni (GPT-4o, gpt-4o-2024-05-13), GPT-4 Turbo (gpt-4-turbo-2024-04-09), LLaVA-NeXT-110B Liu et al.
@@
 The random baseline for the Classification task is calculated by randomly choosing between a “Yes” and a “No” answer for each item, and averaging the results of 100 iterations.
-For ImageBbox, IoU scores above 0.5 are considered a match between model prediction and ground truth.
+For ImageBbox, we follow standard practice in object detection Everingham et al.
+(2010); Lin et al.
+(2014) and consider IoU scores of 0.5 or higher as indicating sufficient overlap between model prediction and ground truth.
 For Naming, ImageLabel, and Explanation, we decide thresholds for good answers by manually examining model answers in various score ranges.
@@
 For Explanation, model answers reaching a ROUGE-1 score of 0.35 and a ROUGE-2 score of 0.087 are considered acceptable.
+Model outputs of different score ranges in these three tasks are presented in Appendix B.
 How well do the models identify humorous multimodal metaphor use?
@@
 Qwen2-VL-7B 0.66 (0.06) 0.36 (0.11) 0.51 (0.03)
-Table 4: Mean F1 scores using 3 different prompts for the Classification task. Qwen2 stands for Qwen2-VL-Instruct. Success rate is always 100%.
-Figure 4: Model performance in multimodal versus monomodal experiments: average F1 score for Classification, sentence similarity score for Naming and ImageLabel, Jaccard index score for CaptionHL, ROUGE-1 for Explanation.
-Success rate is 100% except for Qwen2-VL-7B-Instruct in the monomodal CaptionHL task (97%).
+Table 4: Mean F1 scores using 3 different prompts for the Classification task. Qwen2-VL stands for Qwen2-VL-Instruct. Success rate is always 100%.
 We experiment with other ways to formulate the Classification task, to see whether the high probability of “Yes” answers is associated with the prompt we use.
@@
 The experiment thus proves the reliability of our Classification benchmark results.
+Model Classification Naming ImageBbox Caption
+Pos Neg Avg
+GPT-4 Omni 0.70 0.36 0.53 0.53 (0.23) 0.25 (0.20) 0.50 (0.39)
+Turbo 0.73 0.22 0.48 0.50 (0.23) 0.21 (0.20) 0.33 (0.33)
+LLaVA-NeXT 110B 0.73 0.00 0.37 0.49 (0.19) 0.17 (0.17) 0.62 (0.43)
+8B 0.74 0.01 0.37 0.45 (0.17) 0.15 (0.16) 0.30 (0.31)
+Qwen2-VL 72B 0.73 0.00 0.37 0.47 (0.19) 0.17 (0.19) 0.39 (0.39)
+7B 0.73 0.00 0.36 0.51 (0.20) 0.13 (0.15) 0.38 (0.36)
+Human 0.75 0.71 0.73 0.63 (0.20) 0.73 (0.32) 0.65 (0.41)
+Table 5: Model performance compared with human performance on Classification, Naming, ImageBbox (IoU), and Caption (Jaccard index) tasks.
+Prompts for the models are similar to the annotation guidelines.
+Some of the tests result in success rate less than 100%: $\geq$90% in Naming; $\geq$90% for the Qwen2-VL models and $\geq$80% for the other models in ImageBbox; $\geq$90% for GPT-4 Turbo and Qwen2-VL-72B, 79% for GPT-4o, and $\sim$60% for the other models in CaptionHL.
+Figure 5: Model performance in multimodal versus monomodal experiments: average F1 score for Classification, sentence similarity score for Naming and ImageLabel, Jaccard index score for CaptionHL, ROUGE-1 for Explanation.
+Success rate is 100% except for Qwen2-VL-7B-Instruct in the monomodal CaptionHL task (97%).
+6.3 Comparison with human performance
+Is the models’ compromised performance really due to the their lack of capacity to process humorous multimodal metaphor use, or is it due to the subjectivity of the tasks?
+To answer this question, we design a new prompt that is similar to the annotation guidelines we provided to the second expert annotator—the models are instructed to complete Classification, Naming, ImageBbox, and CaptionHL in one go, following the same steps as the expert annotator.
+We also ask the models to adopt a persona that is representative of the human experts: a linguist who is familiar with CMT and has experience in manual metaphor annotation.
+The prompt is provided in Appendix B.
+It is tested on the smallest model, Qwen2-VL-7B, before being used for the full experiment.
+As shown in Table 5, even when following similar guidelines, none of the models reach the level of agreement between the two expert annotators.
+Model performance is closer to human performance in the Naming task, as compared to the other tasks.
+In CaptionHL, LLaVA-NeXT-110B reaches a Jaccard index score (0.62) that is close to human (0.65), but with a low success rate (62%): for $\sim$40% of the instances the model identifies as metaphorical, it fails to provide a caption annotation.
+In Classification and ImageBbox, the human expert surpasses the models by a large margin.
+Also note that this new approach does not help with the Classification task despite its more detailed instructions about what is metaphor and how to identify humorous metaphor use: most of the models appear to be better at excluding negative cases in the benchmark task (Table 2), although the performance is still much lower than the human expert.
+Nonetheless, the new approach benefits the Naming task and yields higher Jaccard index scores in the CaptionHL task for GPT-4o, LLaVA-NeXT-110B, and Qwen2-VL-7B, albeit with compromised success rates for the two larger models.
+Overall, the new approach of following the manual annotation guidelines is better suited for detailed annotation of humorous metaphor use (Naming, ImageBbox, and CaptionHL).
+As for the Classification task, the models’ inability to identify negative cases seems to go beyond what can be achieved by prompt engineering.
 6.4 Ablation study
@@
 The comparison indicates that the multimodal input data is not adequately utilized by the models.
+Our error analysis provides more support for this.
 6.5 Error Analysis
-Figure 5: Ground truth versus LLaVA-NeXT-110B predictions for the given image-caption pair. Explanations are shortened. For CaptionHL, the model outputs the entire caption as answer.
+Figure 6: Ground truth (left and in the caption) versus LLaVA-NeXT-110B predictions (right) for the given image-caption pair.
+Explanations are shortened.
+For CaptionHL, the model outputs the entire caption as answer.
 Figure 7: All models except GPT-4o predict life is a vacuum for this item in the Naming task.
@@
 We hope this study will encourage more research on MLLMs’ capabilities to process humorous multimodal metaphors.
-Our dataset is annotated by a single expert.
-While this ensures efficient, high-quality, and systematic labelling, it may also introduce a certain level of subjectivity to the dataset.
-We envisage future work exploring alternative interpretations of the image-caption pairs and how MLLMs align with such individuality.
+While we have experimented with promoting model performance through prompt engineering, other possibilities such as few-shot prompting and fine-tuning are beyond the scope of the current study.
 We also acknowledge that the latest reasoning models such as DeepSeek R1 and OpenAI o1 are not included in this study.
@@
 The CapCon corpus has CC-BY-4.0 license.
-Our dataset will be freely accessible on GitHub.
+Our dataset is freely accessible on GitHub.
 Our dataset includes jokes that could be considered offensive, and certain jokes may be inappropriate for a younger audience.
-These data remain in our dataset as they can be valuable for future research.
+These data remain in our dataset as they are a nonremovable part of multimodal humor and can be valuable for future research.
+Appendix A Annotation Guidelines
+(a) Non-metaphorical.
+(b) Metaphorical.
+Figure 8: Image-caption pair that (a) does not involve and (b) involves humorous metaphor use.
+(a) Correct.
+(b) Too small.
+(c) Too big.
+Figure 9: Good and bad examples for bounding box annotation.
+A.1 Objective
+You will annotate New Yorker cartoons and their captions in terms of humorous metaphor use.
+A.2 Labelling Steps
+  1. 1.
+    Look at the cartoon image alone (without referring to the caption). Find any incongruities in the image. You can write these down in the incongruity field, but this is optional.
+  2. 2.
+    Copy and paste the caption to text_hl.
+  3. 3.
+    How does the caption resolve the incongruities in the image?
+If you can answer this question, then you understand the humor of the image-and-caption combination.
+You can write down your understanding in the resolution field, but this is optional.
+    1. (a)
+      If you understand the humor, continue with the annotation;
+    2. (b)
+      Otherwise, select “Discard” for the is_met field, assign “1” to complete and submit your annotation for this item.
+  4. 4.
+    Based on your understanding of the humor, consider whether metaphor use (as detailed in “What counts as metaphor”) is involved in creating that humorous effect.
+Any kind of involvement counts.
+In essence, for each metaphor use you find in the image-and-caption combination, consider whether the humor will still be there when the metaphor use is removed.
+    1. (a)
+      If the answer is an absolute yes, assign “Yes” to is_met and continue with the annotation;
+    2. (b)
+      If the answer is an absolute no, assign “No” to is_met, “1” to complete and submit your annotation;
+    3. (c)
+      If you hesitate, assign “WIDLII” to is_met and continue with the annotation.
+  5. 5.
+    Specify the conceptual metaphor(s) in “X is Y” format.
+“X” should be the target domain; “Y” should be the source domain.
+Separate multiple conceptual metaphors with “; ” (a semicolon followed by a whitespace).
+    1. (a)
+      To decide how general/specific the conceptual metaphor should be, try to relate the instantiation of the metaphor in the image-and-caption combination to other instantiations of the same metaphor (in any modality), and then find a term that can cover them all.
+    2. (b)
+      You can draw inspiration from this (non-exhaustive) list of conceptual metaphors.
+    3. (c)
+      Finishing the image_hl and text_hl annotation first (Step 6) could be helpful if you had difficulty pinpointing the two domains.
+  6. 6.
+    Annotate the metaphor-related image areas (image_hl) and text fragments (text_hl):
+    1. (a)
+      To annotate metaphor-related image areas, try to assign each object/element of the image to either the target or the source domain of the conceptual metaphors you have annotated.
+Typically, the majority of the image should belong to the same domain (e.g., the target domain); that one object/element that belongs to the other domain (e.g., the source domain) is the metaphor-related image area.
+Draw a image_hl bounding box for each such “odd” object/element (e.g., if the minority domain is CAT and there are 3 cats in the image, draw 3 bounding boxes, one box for each cat);
+    2. (b)
+      Do the same for the caption.
+In the text_hl field, where you already have the caption copied and pasted, surround each word/phrase/clause that belongs to the “odd” domain with a pair of <i></i> tag.
+If the entire caption really belongs to a single domain, surround the entire caption in a pair of <i></i> tag.
+Note that image_hl and text_hl do not have to belong to the same domain.
+A.3 Mandatory fields
+  1. 1.
+    Every item should have a is_met value.
+  2. 2.
+    If is_met value is “Yes” or “WIDLII”, the following fields should also be filled: image_hl, text_hl (with <i>highlighted text</i>), and conceptual_metaphor.
+A.4 What counts as metaphor
+Following the Conceptual Metaphor Theory, we define metaphors as conceptual mappings between two difference domains. Therefore:
+  1. 1.
+    Personification and zoomorphism are considered metaphor use. Personification is metaphor use with HUMAN as the source domain; zoomorphism is metaphor use with ANIMAL as the source domain.
+  2. 2.
+    Metonymy within the same conceptual domain does not count as metaphor use.
+  3. 3.
+    Idioms are not metaphors, unless the underlying cross-domain mapping is strictly required to make sense of the humor.
+  4. 4.
+    Puns can indicate metaphor use. Whether they are metaphorical depends on whether the two meanings can be attributed to some sort of cross-domain mapping.
+A.5 Non-metaphorical examples
+Example 1: Figure 8(a).
+Incongruity: Giant beach ball.
+Resolution: Whoever that was confident to catch the ball is now under it.
+Why not metaphorical: It’s literally a giant ball; someone’s literally under it in the pool.
+Example 2 and 3 are omitted due to space limit.
+A.6 Metaphorical examples
+Example 1: Figure 8(b).
+Incongruity: Bear in a restaurant.
+Resolution: The bear is a waiter.
+Why metaphorical: Restaurant waiters are human beings.
+Conceptual metaphor: animals are human workers.
+Image annotation: See Figure 9(a).
+Caption annotation: “Thanks for checking on the <i>salmon</i>.
+How’s the lamb?”.
+Example 2 and 3 are omitted due to space limit.
+A.7 Bounding box examples
+The bounding box should fit tightly around the object of interest—for example, the bear in Figure 9(a).
+Make sure the entire object is included; Figure 9(b) is therefore incorrect.
+But also avoid including excessive empty space around the object, such as Figure 9(c).
+Appendix B Additional prompts and model outputs
+Model Output vs. Ground Truth Score
+Dogs are people 0.83
+vs. Animals are humans
+Temperature is music 0.73
+vs. Temperature is pitch
+Thoughts are bathwater 0.63
+vs. Psychotherapy is a bath
+Parenting is egg incubation 0.60
+vs. Human baby is an egg
+Humor is a drug 0.55
+vs. Preaching is a joke
+Faith is a wedding 0.45
+vs. Alcohol is god
+Work is a tool 0.35
+vs. Psychotherapy is a bath
+Slaying a dragon is a task 0.26
+vs. Modern man is knight
+Ending a relationship is falling off a cliff vs. Social media is physical world 0.17
+Table 6: Sample model outputs in the Naming benchmark task and their cosine similarity scores compared with ground truth.
+Model Output Ground Truth Score
+Gun Pistol 0.86
+Teddy-bear Toy bear 0.75
+Meteor Asteroid 0.69
+Tree stump Trunk 0.62
+Ball Egg 0.57
+Hat Cellphone 0.47
+Alligator Customer 0.40
+Laptop Asteroid 0.32
+Beer Human-shaped hole 0.21
+Broom Alien spaceship 0.10
+Table 7: Sample model outputs in the ImageLabel benchmark task and their cosine similarity scores compared with ground truth.
+Figure 10: Sample model outputs in the Explanation benchmark task and their ROUGE-1/ROUGE-2 scores as compared to ground truth.
+Figure 11: Prompt for comparing model and human performance.
+Table 6, Table 7, and Figure 10 provide example model outputs in the ImageLabel, Naming, and Explanantion benchmark tasks respectively.
+The prompt used in Section 6.3 to compare model and human performance is provided in Figure 11.